# Physics-inspired Generative Models in Medical Imaging: A Review

### v1

@@ 2.5.2 Trade-off robustness for rigidity by varying $D$
In PFGM++, $D$ is treated as a hyperparameter to be tuned. In particular, Xu et al. [109] demonstrates that by tuning $D\in(0,\infty)$ one can trade off robustness (insensitive to missteps) for rigidity (effective learning). In particular, as first shown for PFGM [108], PFGM++ is less susceptible to estimation errors than diffusion. Xu et al. [109] demonstrates the superior robustness of PFGM++ to three sources of errors: noise injection, large step-size, and post-training quantization. At training time, for large $D$ and $r$, the marginal distribution $p_{r}(\bm{x})$ is supported on a sphere with radius $r\sqrt{N/D}.$ If the sample trajectory deviates from this $r$-norm relation in the training data, the backward ODE can lead to unexpected results. Hence, setting $D$ small, moving away from diffusion models, increase robustness. However, a small $D$ results in a heavy-tailed input distribution and a broad input range is challenging to handle for any finite capacity network. Thus, optimal performance occurs for some $D\in(0,\infty)$ that requires tuning for the problem at hand.
2.6 Latent Diffusion Models (LDM)
@@
Latent Diffusion Models (LDM) [71] innovate by operating in a latent space—a compressed representation of the data—rather than directly in the data’s original space. This method significantly reduces computational costs while maintaining, or even enhancing, the quality of generated samples. In particular, Rombach et al. [71] suggest a perceptual image compression via an autoencoder trained with a combination of perceptual and adversarial objectives. For a RGB image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ the encoder $\mathcal{E}$ encodes $\bm{x}$ into the latent representation $\bm{z}=\mathcal{E}(\bm{x})\in\mathbb{R}^{h\times w\times c}.$ The decoder $\mathcal{D}$ reverse this process, reconstructing the image from its latent representation $\tilde{\bm{x}}=\mathcal{D}(\mathcal{E}(\bm{x})).$ Rombach et al. [71] investigates a range of down-sampling factors $f=H/h=W/h$ and finds that $f=\{4,8\}$ yields a good trade-off between computational complexity and image fidelity. In addition, Rombach et al. [71] introduces a very general conditioning mechanism to model conditional distributions $p(\bm{z}|\bm{y})$ by augmenting the diffusion model backbone with a cross-attention mechanism [93], allowing the network to efficiently learn embeddings from varying modalities via domain specific encoders $\tau_{\bm{\theta}}(y)\in\mathbb{R}^{M\times d_{\tau}}.$ A schematic overview of LDM is available in Figure 2 (c).
@@
LDM are celebrated for their ability to produce high-quality outputs with reduced computational demands. By leveraging the efficiencies of latent space, LDM offer a powerful solution for generating complex data, such as high-resolution images or intricate text sequences, making them a valuable tool in the arsenal of generative modeling techniques. A particular instance of LDM, a text-to-image diffusion model called Stable Diffusion, https://github.com/CompVis/stable-diffusion, has gained unprecedented popularity for its effectiveness in image generation and manipulation.
@@ 2.7 Accelerated Sampling Schemes
The iterative nature of physics-inspired deep generative models is a double-edge sword. On the one hand, it allows for a flexible trade-off between compute and image quality, in addition to zero-shot inverse problem solving. On the other hand, this means that the number of function evaluations (NFE) required may be order of magnitudes larger than for single-step samplers. This computationally heavy sampling limit their usage in applications where speed is of the essence.
One line of work to accelerate sampling involves developing more efficient samplers. In particular, a salient trend is to move from stochastic to deterministic samplers. This was done for DDPM in Denoising Diffusion Implicit Models (DDIM) [80] and for score-based diffusion models in Karras et al. [52]. Notably, relatively newer methods such as PFGM [108] and PFGM++ [109] also follow this trend, focusing directly on the ODE instead of the SDE.
Another prominent development in the quest for more efficient sampling are consistency models [85, 81]. Consistency models are a class of generative models that build on the continuous-time formulation of diffusion models [84, 52] to enable single-step, $\text{NFE}=1$, sampling. These models achieve this by directly mapping noise to data, enabling high-quality sample generation in a single step, by enforcing self consistency at training time. An illustration of the mapping learned in consistency models is available in Figure 2 (d). Let $\bm{x}_{\sigma}$ denote an intermediate step on the solution trajectory $\{\bm{x}_{\sigma}\}_{\sigma\in[\sigma_{\text{min}},\sigma_{\text{max}}]}$, where $\bm{x}_{\sigma_{\text{min}}}\sim p(\bm{x})$ and $\bm{x}_{\sigma_{\text{max}}}\sim\mathcal{N}(\bm{0},\sigma^{2}_{\text{max}}).$ Consistency models then strive to learn a consistency function
@@
where $c_{\text{skip}}(\sigma)$ and $c_{\text{out}}(\sigma)$ are differentiable functions such that $c_{\text{skip}}(\epsilon)=1$ and $c_{\text{out}}(\epsilon)=0$ and $f_{\bm{\theta}}(\bm{x},\sigma)$ denotes our approximated consistency function via the deep neural network $F_{\bm{\theta}}(\bm{x},t).$ The domain $[\sigma_{\text{min}},\sigma_{\text{max}}]$ is discretized using a sequence of noise scales $\sigma_{\text{min}}=\sigma_{1}<\sigma_{2}<...<\sigma_{N}=\sigma_{\text{max}}.$ Following Karras et al. [52]Song et al. [85] sets $\sigma_{i}=(\sigma_{\text{min}}^{1/\rho}+\frac{i-1}{N-1}(\sigma_{\text{max}}^{1/\rho}-\sigma_{\text{min}}^{1/\rho}))^{\rho}$ with $\rho=7$ for $i=1,...,N.$
@@
where $0\leq\mu\leq 1$ is the decay rate. Since $\nabla_{\bm{x}}\log p_{\sigma_{i+1}}(\bm{x})$ is not known, minimizing the objective in Equation 20 is not feasible. Consistency training and consistency distillation are two algorithms to address this issue [85]. Both approaches use an approximation to generate a pair of adjacent points on the solution trajectory, approximately adhering to Equation 20 Consistency distillation, as the name implies, distills a pre-trained diffusion model into a consistency model and simply replacing the score-function in Equation 20 by a pre-trained score network. Consistency training, on the other hand, is a stand-alone model and estimates the two adjacent points on the solution trajectory via the perturbation kernel. In particular, consistency training obtains the pair of adjacent points on the same solution trajectory $(\bm{x}_{\sigma_{i+1}},\breve{\bm{x}}_{\sigma_{i}})$ via $\bm{x}_{\sigma_{i+1}}=\bm{x}+\sigma_{i+1}\bm{z}$ and $\breve{\bm{x}}_{\sigma_{i}}=\bm{x}+\sigma_{i}\bm{z}$ using the same $\bm{x}$ and $\bm{z}.$ Consistency training then employ the loss
@@
Using consistency models to achieve high fidelity output but with significantly less compute has been extended to the case of LDM in Luo et al. [66], with a method called Latent Consistency Distillation (LCD). In addition to generalizing the consistency matching framework to text-to-image LDM, the authors also generalizes the idea of enforcing consistency between adjacent time steps $t_{n+1}$ and $t_{n}$ to consistency between time steps that are $k$ steps apart, that is $t_{n+k}$ and $t_{n}.$ This methodology generates impressive state of the art few-step, $\text{NFE}<8$, text-to-image generation.
Consistency trajectory models (CTM) [55] learn to estimate both the integrand (the score function) and the integral over any time horizon of the PF ODE, providing a unifying framework which includes consistency models and diffusion models as special cases. This setup enables anytime-to-anytime jumps, which is leveraged efficiently to tradeoff compute for image quality via novel sample techniques. In particular, CTM achieves impressive state of the art performance for unconditional image generation on CIFAR-10 and ImageNet at $64\times 64$ resolution datasets with $\text{NFE}=1$. Notably, however, CTM is using a combination of a distillation loss, DSM loss, and an adversarial loss. The inclusion of an adversarial loss can lead to training instability, tentatively negating a key advantage of diffusion models over GANs.
@@ 3.1 Image Reconstruction
Deep learning methods for CT image reconstruction have commonly employed supervised learning to map partial measurements to complete images [98, 67], relying on extensive datasets of paired CT images and measurements. However, variations in the measurement process, such as changing CT projections, require new data collection and model retraining. This limits generalization and can sometimes lead to counter-intuitive instabilities [5]. Recently, diffusion models have addressed these issues by introducing unsupervised training, which provides effective priors without the need for labeled data [83, 21]. These models, which train on high-quality images/data to fill in gaps in down-sampled data, eliminating the need for matched datasets, supporting unsupervised learning and improving tomographic reconstruction quality. This sub-section delves into the application of diffusion models in tomographic reconstruction, emphasizing their role in producing superior images and diagnostic performance.
@@ 3.1.1 CT Reconstruction
The initial work reported in [83] adapted the diffusion model to medical image reconstruction. This score-based generative model involves solving inverse problems by generating samples from the posterior distribution $p({\mathbf{x}}\mid{\mathbf{y}})$. This is achieved by conditioning a stochastic process $\{{\mathbf{x}}_{t}\}_{t\in[0,1]}$ on an observation ${\mathbf{y}}$, leading to the conditional process $\{{\mathbf{x}}_{t}\mid{\mathbf{y}}\}_{t\in[0,1]}$. The core of this approach is the conditional stochastic process as a reverse-time stochastic differential equation (SDE):
@@
Along this direction, Cong et al. [26] improved the diffusion model-based tomographic reconstruction approach to perform the maximum a posteriori (MAP) reconstruction in the Bayesian framework (cf. the introduction) utilizing a Gaussian mixture model that more accurately captures discrepancies between the distributions of reconstructed images and real counterparts, respectively. This integration facilitates the derivation of a novel score-matching formula implemented via deep learning. The performance of this Gaussian mixture score-matching method was evaluated on both public medical image datasets and clinical raw CT datasets, demonstrating its effectiveness and robustness compared with traditional methods [89] and existing diffusion models.
@@
To alleviate the computational expense of diffusion models, Liu et al. [62] introduced a Fourier diffusion approach, adapted from [91], allowing for many fewer time steps than the standard scalar diffusion model. This technique models the end-to-end process of X-ray transmission tomography and FBP to estimate image properties in sparse CT systems. They demonstrated the effectiveness of their method in reducing computational requirements while maintaining image quality on a simulated breast cone-beam CT system using sparse view acquisition.
@@
To further alleviate the computational bottleneck in DDPMs, the patch-based DDPM, [103] was designed so that the diffusion-based inverse problem can be solved in a distributed parallel fashion. This novel approach was demonstrated for sparse-view CT reconstruction. The network, trained on patches from fully sampled projection data, enables unsupervised inpainting of down-sampled data without the need for paired datasets. By employing a patch-based approach, this method allows for distributed parallel processing, mitigating memory issues. Additionally, a projection-domain patch-based DDPM method is inspired by using a U-Net [72] to learn the reverse diffusion process on fully sampled projection patches. The reconstructed patches are finally assembled to form a high-quality projection dataset for FBP image reconstruction, Such an integration of unsupervised learning and parallel computing capability makes the diffusion model suitable for large-scale tasks. Indeed, Xia et al. [104] further refined the approach for breast CT reconstruction, which uses parallel DDPM in both image and sinogram domains for sparse-view reconstruction, significantly reducing radiation doses while maintaining competitive results.
Using Poisson Flow Consistency Models (PFGM), Ge et al. [34] introduced the Joint Condition and Circle-Supervision algorithm for multiphase contrast-enhanced CT (CECT), optimizing image quality across various imaging phases. This strategy integrates temporal progression and spatial correlations to reduce the involved radiation dose.
Recently, Hein et al. [44] presented a PFGM++ approach to photon-counting CT (PCCT) image denoising via posterior sampling. Their strategy employs a consistency-enforced sampling scheme, using a ‘hijacking’ technique that integrates noisy observations directly into the reverse generative process [79, 23, 64]. This avoids complex regularization, effectively approximating the true image from noisy data [53].  Hein et al. [45] further developed this framework with the Posterior Sampling Poisson Flow Consistency Model (PS-PFCM), which optimizes computational efficiency, leveraging a “consistency distillation" process and enabling single-step or few-step function evaluations [86].
@@ 3.1.2 MRI Reconstruction
Similar to deep CT image reconstruction, deep MRI image reconstruction can be viewed in two categories: model-driven and data-driven methods [36] but no clear boundary between these methods. Model-driven methods rely on mathematically derived optimization equations and incorporate artificial neural networks (ANNs) to enhance accuracy and efficiency [75]. Data-driven methods extract patterns and correlations from large datasets using advanced deep learning techniques in an end-to-end fashion [7]. Both methods can improve MRI reconstruction, balancing accuracy, efficiency, and robustness in different ways.
Chung et al. [24] combine the Krylov subspace method [61] with the diffusion model for complex imaging tasks, such as multi-coil MRI and 3D CT scans. This method, called Decomposed Diffusion Sampling (DDS), uses the fast-converging traits of the Krylov subspace method, known from traditional optimization studies, together with the strong generation abilities of the contemporary diffusion models. The key to DDS is its use of the manifold constrained gradient (MCG) [22]. DDS shows that the tangent space at a cleaned-up sample, according to Tweedie’s formula, matches a Krylov subspace. This insight allows the method to perform multiple conjugate gradient (CG) updates within this space, simplifying the process significantly. Mathematically, this process is expressed as
@@
Chung et al. [25] introduced DiffusionMBIR, a novel method that combines model-based iterative reconstruction (MBIR) with 2D diffusion modeling to meet the 3D image reconstruction challenges like sparse-view tomography and compressed sensing MRI. This approach marries the precision of MBIR with the generative power of the diffusion model, enabling efficient, high-fidelity reconstructions on a single commodity GPU. The innovation of DiffusionMBIR lies in augmenting a 2D diffusion prior with a model-based total variation (TV) prior, applied specifically in the longitudinal direction at test time. This method uses the alternating direction method of multipliers (ADMM) [10] to integrate measurement information and enforce consistency across slices:
@@
To further enhance MRI quality, Geng et al. [36] introduced a comprehensive detail-preserving method using multiple diffusion models to extract structural and textural features in the k-space. Traditional diffusion models often struggle with capturing intricate details accurately [74, 40]. The proposed method employs virtual binary modal masks to refine the value range of k-space data, which places the model focus within a highly adaptive center window. An inverted pyramid structure is used, allowing for a multi-scale representation. Empirically, a step-by-step refinement significantly enhances detailed image features on the selected benchmark datasets.

### v2

@@ 2.5.2 Trade-off robustness for rigidity by varying $D$
In PFGM++, $D$ is treated as a hyperparameter to be tuned. In particular, Xu et al. [114] demonstrate that by tuning $D\in(0,\infty)$ one can trade off robustness (insensitive to missteps) for rigidity (effective learning). In particular, as first shown for PFGM [113], PFGM++ are less susceptible to estimation errors than SDMs. Xu et al. [114] demonstrate the superior robustness of PFGM++ to three sources of errors: noise injection, large step-size, and post-training quantization. At training time, for large $D$ and $r$, the marginal distribution $p_{r}(\bm{x})$ is supported on a sphere with radius $r\sqrt{N/D}.$ If the sample trajectory deviates from this $r$-norm relation in the training data, the backward ODE can lead to unexpected results. Hence, setting $D$ small, moving away from SDMs, increase robustness. However, a small $D$ results in a heavy-tailed input distribution and a broad input range is challenging to handle for any finite capacity network. Thus, optimal performance occurs for some $D\in(0,\infty)$ that requires tuning for the problem at hand.
2.6 Latent DMs (LDMs)
@@
LDMs [76] innovate by operating in a latent space—a compressed representation of the data—rather than directly in the data’s original space. This method significantly reduces computational costs while maintaining, or even enhancing, the quality of generated samples. In particular, [76] suggest a perceptual image compression via an autoencoder trained with a combination of perceptual and adversarial objectives. For a RGB image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ the encoder $\mathcal{E}$ encodes $\bm{x}$ into the latent representation $\bm{z}=\mathcal{E}(\bm{x})\in\mathbb{R}^{h\times w\times c}.$ The decoder $\mathcal{D}$ reverse this process, reconstructing the image from its latent representation $\tilde{\bm{x}}=\mathcal{D}(\mathcal{E}(\bm{x})).$ Rombach et al. [76] investigate a range of down-sampling factors $f=H/h=W/h$ and finds that $f=\{4,8\}$ yields a good trade-off between computational complexity and image fidelity. In addition, Rombach et al. [76] introduce a very general conditioning mechanism to model conditional distributions $p(\bm{z}|\bm{y})$ by augmenting the DM backbone with a cross-attention mechanism [98], allowing the network to efficiently learn embeddings from varying modalities via domain specific encoders $\tau_{\bm{\theta}}(y)\in\mathbb{R}^{M\times d_{\tau}}.$ A schematic overview of LDMs is available in Figure 2 (c).
@@
LDMs are celebrated for their ability to produce high-quality outputs with reduced computational demands. By leveraging the efficiencies of latent space, LDMs offer a powerful solution for generating complex data, such as high-resolution images or intricate text sequences, making them a valuable tool in the arsenal of generative modeling techniques. A particular instance of LDMs, a text-to-image DM called Stable Diffusion ( https://github.com/CompVis/stable-diffusion) has gained unprecedented popularity for its effectiveness in image generation and manipulation.
@@ 2.7 Accelerated Sampling Schemes
The iterative nature of physics-inspired GMs is a double-edge sword. On the one hand, it allows for a flexible trade-off between compute and image quality, in addition to zero-shot inverse problem solving. On the other hand, this means that the number of function evaluations (NFE) required may be order of magnitudes larger than for single-step samplers. This computationally heavy sampling limit their usage in applications where speed is of the essence.
One line of work to accelerate sampling involves developing more efficient samplers. In particular, a salient trend is to move from stochastic to deterministic samplers. This was done for DDPMs in Denoising Diffusion Implicit Models (DDIMs) [85] and for SDMs in Karras et al. [54]. Notably, relatively newer methods such as PFGMs [113] and PFGM++ [114] also follow this trend, focusing directly on the ODE instead of the SDE.
Another prominent development in the quest for more efficient sampling are consistency models [90, 86]. Consistency models are a class of GMs that build on the continuous-time formulation of DMs [89, 54] to enable single-step, $\text{NFE}=1$, sampling. These models achieve this by directly mapping noise to data, enabling high-quality sample generation in a single step, by enforcing self consistency at training time. An illustration of the mapping learned in consistency models is available in Figure 2 (d). Let $\bm{x}_{\sigma}$ denote an intermediate step on the solution trajectory $\{\bm{x}_{\sigma}\}_{\sigma\in[\sigma_{\text{min}},\sigma_{\text{max}}]}$, where $\bm{x}_{\sigma_{\text{min}}}\sim p(\bm{x})$ and $\bm{x}_{\sigma_{\text{max}}}\sim\mathcal{N}(\bm{0},\sigma^{2}_{\text{max}}).$ Consistency models then strive to learn a consistency function
@@
where $c_{\text{skip}}(\sigma)$ and $c_{\text{out}}(\sigma)$ are differentiable functions such that $c_{\text{skip}}(\epsilon)=1$ and $c_{\text{out}}(\epsilon)=0$ and $f_{\bm{\theta}}(\bm{x},\sigma)$ denotes our approximated consistency function via the deep neural network $F_{\bm{\theta}}(\bm{x},t).$ The domain $[\sigma_{\text{min}},\sigma_{\text{max}}]$ is discretized using a sequence of noise scales $\sigma_{\text{min}}=\sigma_{1}<\sigma_{2}<...<\sigma_{N}=\sigma_{\text{max}}.$ Following Karras et al. [54]Song et al. [90] set $\sigma_{i}=(\sigma_{\text{min}}^{1/\rho}+\frac{i-1}{N-1}(\sigma_{\text{max}}^{1/\rho}-\sigma_{\text{min}}^{1/\rho}))^{\rho}$ with $\rho=7$ for $i=1,...,N.$
@@
where $0\leq\mu\leq 1$ is the decay rate. Since $\nabla_{\bm{x}}\log p_{\sigma_{i+1}}(\bm{x})$ is not known, minimizing the objective in Equation 19 is not feasible. Consistency training and consistency distillation are two algorithms to address this issue [90]. Both approaches use an approximation to generate a pair of adjacent points on the solution trajectory, approximately adhering to Equation 20. Consistency distillation, as the name implies, distills a pre-trained DM into a consistency model and simply replacing the score-function in Equation 20 by a pre-trained score network. Consistency training, on the other hand, is a stand-alone model and estimates the two adjacent points on the solution trajectory via the perturbation kernel. In particular, consistency training obtains the pair of adjacent points on the same solution trajectory $(\bm{x}_{\sigma_{i+1}},\breve{\bm{x}}_{\sigma_{i}})$ via $\bm{x}_{\sigma_{i+1}}=\bm{x}+\sigma_{i+1}\bm{z}$ and $\breve{\bm{x}}_{\sigma_{i}}=\bm{x}+\sigma_{i}\bm{z}$ using the same $\bm{x}$ and $\bm{z}.$ Consistency training then employ the loss
@@
Using self-consistency to achieve high fidelity output with significantly less compute has been extended to the case of LDM in Luo et al. [71], with a method called Latent Consistency Distillation (LCD). In addition to generalizing the consistency matching framework to text-to-image LDM, the authors also generalizes the idea of enforcing consistency between adjacent time steps $t_{n+1}$ and $t_{n}$ to consistency between time steps that are $k$ steps apart, that is $t_{n+k}$ and $t_{n}.$ This methodology generates impressive state of the art few-step, $\text{NFE}<8$, text-to-image generation.
Consistency trajectory models (CTMs) [57] learn to estimate both the integrand (the score function) and the integral over any time horizon of the PF ODE, providing a unifying framework which includes consistency models and DMs as special cases. This setup enables anytime-to-anytime jumps, which is leveraged efficiently to trade off compute for image quality via novel sample techniques. In particular, CTMs achieve impressive state of the art performance for unconditional image generation on CIFAR-10 and ImageNet at $64\times 64$ resolution datasets with $\text{NFE}=1$. Notably, however, CTMs are using a combination of a distillation loss, DSM loss, and an adversarial loss. The inclusion of an adversarial loss can lead to training instability, tentatively negating a key advantage of DMs over GANs.
@@ 3.1 Image Reconstruction
Deep learning methods for CT image reconstruction have commonly employed supervised learning to map partial measurements to complete images [103, 72], relying on extensive datasets of paired CT images and measurements. However, variations in the measurement process, such as changing CT projections, require new data collection and model retraining. This limits generalization and can sometimes lead to counter-intuitive instabilities [5]. Recently, DMs have addressed these issues by introducing unsupervised training, which provides effective priors without the need for labeled data [88, 22]. These models, which train on high-quality images/data to fill in gaps in down-sampled data, eliminating the need for matched datasets, supporting unsupervised learning and improving tomographic reconstruction quality. The following sub-sections delve into the application of DMs in tomographic, MRI, and PET reconstruction, emphasizing their role in producing superior images and diagnostic performance.
@@ 3.1.1 CT Reconstruction
The initial work reported in [88] adapted the diffusion model to medical image reconstruction. This score-based generative model, as outlined in Section 2.3, involves solving inverse problems by generating samples from the posterior distribution $p({\mathbf{x}}\mid{\mathbf{y}})$. This is achieved by conditioning a stochastic process $\{{\mathbf{x}}_{t}\}_{t\in[0,1]}$ on an observation ${\mathbf{y}}$, leading to the conditional process $\{{\mathbf{x}}_{t}\mid{\mathbf{y}}\}_{t\in[0,1]}$. The core of this approach is the conditional stochastic process as a reverse-time SDE:
@@
Along this direction, Cong et al. [27] improved the diffusion model-based tomographic reconstruction approach to perform the MAP reconstruction in the Bayesian framework utilizing a Gaussian mixture model that more accurately captures discrepancies between the distributions of reconstructed images and real counterparts, respectively. This integration facilitates the derivation of a novel score-matching formula implemented via deep learning. The performance of this Gaussian mixture score-matching method was evaluated on both public medical image datasets and clinical raw CT datasets, demonstrating its effectiveness and robustness compared with traditional methods [94] and existing DMs.
@@
As an example of accelerating techniques in DMs for alleviating computational burdens, discussed in Section 2.7,  Liu et al. [66] have adopted a Fourier diffusion approach originally from [96]. This method considerably cuts down the number of required time steps compared to traditional scalar DMs. Specifically tailored for X-ray transmission tomography and FBP processes, their technique efficiently estimates image properties in sparse CT systems. They validated the efficacy of this approach by demonstrating its ability to maintain image quality while reducing computational demands in a simulated breast cone-beam CT system with sparse view acquisition.
@@
To further alleviate the computational bottleneck in DDPMs, as investigated in Section 2.1, the patch-based DDPM, [108] was designed so that the diffusion-based inverse problem can be solved in a distributed parallel fashion. This novel approach was demonstrated for sparse-view CT reconstruction. The network, trained on patches from fully sampled projection data, enables unsupervised inpainting of down-sampled data without the need for paired datasets. By employing a patch-based approach, this method allows for distributed parallel processing, mitigating memory issues. Additionally, a projection-domain patch-based DDPM method is inspired by using a U-Net [77] to learn the reverse diffusion process on fully sampled projection patches. The reconstructed patches are finally assembled to form a high-quality projection dataset for FBP image reconstruction, Such an integration of unsupervised learning and parallel computing capability makes the diffusion model suitable for large-scale tasks. Indeed, Xia et al. [109] further refined the approach for breast CT reconstruction, which uses parallel DDPM in both image and sinogram domains for sparse-view reconstruction, significantly reducing radiation doses while maintaining competitive results.
Using PFGM, as detailed in Section 2.4Ge et al. [35] introduced the Joint Condition and Circle-Supervision algorithm for multiphase contrast-enhanced CT (CECT), optimizing image quality across various imaging phases. This strategy integrates temporal progression and spatial correlations to reduce the involved radiation dose. To further enhancement of this approach, as outlined in Section 2.5, Hein et al. [46] presented a PFGM++ approach to photon-counting CT (PCCT) image denoising via posterior sampling. Their strategy employs a consistency-enforced sampling scheme, using a ‘hijacking’ technique that integrates noisy observations directly into the reverse generative process [84, 24, 69]. This avoids complex regularization, effectively approximating the true image from noisy data [55].  Hein et al. [47] further developed this framework with the Posterior Sampling Poisson Flow Consistency Model (PS-PFCM), which optimizes computational efficiency, leveraging a “consistency distillation" process and enabling single-step or few-step function evaluations [91].
@@ 3.1.2 MRI Reconstruction
Similar to deep CT image reconstruction, deep MRI image reconstruction can be viewed in two categories: model-driven and data-driven methods [37], although there are not clear boundary between these methods. Model-driven methods rely on mathematically derived optimization equations and incorporate artificial neural networks (ANNs) to enhance accuracy and efficiency [80]. Data-driven methods extract patterns and correlations from large datasets using advanced deep learning techniques in an end-to-end fashion [7]. Both methods can improve MRI reconstruction, balancing accuracy, efficiency, and robustness in different ways.
Chung et al. [25] combine the Krylov subspace method [64] with the diffusion model for complex imaging tasks, such as multi-coil MRI and 3D CT scans. This method, called Decomposed Diffusion Sampling (DDS), uses the fast-converging traits of the Krylov subspace method, known from traditional optimization studies, together with the strong generation abilities of the contemporary DMs. The key to DDS is its use of the manifold constrained gradient (MCG) [23]. DDS shows that the tangent space at a cleaned-up sample, according to Tweedie’s formula, matches a Krylov subspace. This insight allows the method to perform multiple conjugate gradient (CG) updates within this space, simplifying the process significantly. Mathematically, this process is expressed as
@@
Chung et al. [26] introduced DiffusionMBIR, a novel method that combines model-based iterative reconstruction (MBIR) with 2D diffusion modeling to meet the 3D image reconstruction challenges like sparse-view tomography and compressed sensing MRI. This approach marries the precision of MBIR with the generative power of the diffusion model, enabling efficient, high-fidelity reconstructions on a single commodity GPU. The innovation of DiffusionMBIR lies in augmenting a 2D diffusion prior with a model-based TV prior, applied specifically in the longitudinal direction at test time. This method uses the alternating direction method of multipliers (ADMM) [11] to integrate measurement information and enforce consistency across slices:
@@
To further enhance MRI quality, Geng et al. [37] introduced a comprehensive detail-preserving method using multiple DMs to extract structural and textural features in the k-space. Traditional DMs often struggle with capturing intricate details accurately [79, 42]. The proposed method employs virtual binary modal masks to refine the value range of k-space data, which places the model focus within a highly adaptive center window. An inverted pyramid structure is used, allowing for a multi-scale representation. Empirically, a step-by-step refinement significantly enhances detailed image features on the selected benchmark datasets.